OpenStudy (anonymous):

A local hamburger shop sold a combined total of 439 hamburgers and cheeseburgers on Monday. There were 61 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Monday?

OpenStudy (anonymous):

C + ( C + 61) = 439 2C= 439 - 61 =>C = 378/2 = 189 cheese => H = 439 - 189 = 250 Hamburgers